Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 25     Entry time: 19 Feb 2025, 21:25
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery
-12.4 keV, 202 pulses, 0.5 MHz, 1mJ/pulse

Optical laser delivery
- ns laser (660 nm wavelength) for jet illumination

Achievements/Observations
- NKB focus optimized
- One PCS user program successfully finished
- Second PCS user program was just started

Issues
- Right before we could start with beam defocusing, the beam was gone around 8:45 pm. BKR called to let us know that there will be no beam today (cryo issues)

ELOG V3.1.4-7c3fd00